OpenStudy (anonymous):

If I am doing the derivative of cross product of two vectors - can I first take the cross product then take the derivative - or do I have to use the cross product rule for derivatives

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

I am pretty sure it will work both ways just one longer than the other

OpenStudy (michele_laino):

we can show that the first derivative of a cross product between two vectors, say a, and b is given by the subsequent identity: \[\Large \frac{d}{{dt}}\left( {{\mathbf{a}} \times {\mathbf{b}}} \right) = \frac{{d{\mathbf{a}}}}{{dt}} \times {\mathbf{b}} + {\mathbf{a}} \times \frac{{d{\mathbf{b}}}}{{dt}}\]

OpenStudy (jhannybean):

Known as the Product Rule

OpenStudy (jhannybean):

Both methods should work out fine, like you said, one would just take a little longer.